OpenStudy (anonymous):

Solve and express the solution set in simplest form. 6x-1/5=3/1

OpenStudy (texaschic101):

6x - 1/5 = 3/1 If your like me and don't like to deal with fractions, multiply the entire equation by the LCD, which is 5. Giving us : 30x - 1 = 15 now add 1 to both sides 30x = 15 + 1 30x = 16 x = 16/30 which reduces to 8/15 or with fractions... 6x - 1/5 = 3/1 (add 1/5 to each side) 6x = 3/1 + 1/5 6x = 15/5 + 1/5 6x = 16/5 x = (16/5)/6 x = 16/5 x 1/6 x = 16/30 which reduces to 8/15 either way, you will get the same answer. Any questions ?